OpenStudy (anonymous):

I've tried to solve x^2-3y^2=-11 and x-y=-1 multiple times, but I always get different answers. Could someone help me with this and explain what method I need to use?

OpenStudy (anonymous):

I would go with substitution method

OpenStudy (anonymous):

x=y-1 (y-1)^2 +3y^2=-11

OpenStudy (anonymous):

y^2-2y+1 +3y^2=-11

OpenStudy (anonymous):

no real solution for y

OpenStudy (anonymous):

The solutions for the following equations\[{x^2 - 3 y^2 = -11, x - y = -1}\]using Mathematica is\[\{\{x\to -4,y\to -3\},\{x\to 1,y\to 2\}\} \]